The area of a rectangle is 30cm squared. The width of the rectangle is x cm and the rectangle is 6cm longer than it is wide. Form an equation in x and simplify it.
Answer by richwmiller(17219)

x*(x+6)=30
x^2+6x-30=0
x=3.245
x+6=9.245